OpenStudy (anonymous):

What is the slope of the line passing through the points (0, 4) and (−8, −1)?

OpenStudy (anonymous):

I got 5/12 for an answer

OpenStudy (alexandervonhumboldt2):

use slope formula: \[\frac{ y_1-y_2 }{ x_1-x_2}\]

OpenStudy (alexandervonhumboldt2):

\[\frac{ 4-1 }{ 0+8 }=5/8\]

OpenStudy (anonymous):

ohh ok i got a little counfused thx

OpenStudy (alexandervonhumboldt2):

lol :)